# Iterated sampling importance resampling with adaptive number of proposals

### v1

@@ 6. Convexity and monotonicity of i-SIR covariances and acceptance rates
We will now give the definitions we require. We denote the natural numbers as $\mathbb{N}=\{0,1,2,\dots\}$ and positive natural numbers as $\mathbb{N}_{+}=\{1,2,\dots\}$. Throughout the upcoming sections we have fixed $\pi$ and $q$ as mentioned in Section 1 (more detailed in Appendix B). We use the shorthand notation $\pi(\mathrm{d}x)=\pi(x)\mathrm{d}\mu(x)$ and $q(\mathrm{d}x)=q(x)\mathrm{d}\mu(x)$, where $\mu$ is the dominating measure. When we speak of atoms, we mean atoms with respect to $\pi$. The support of $\pi$ is denoted by $\mathbb{S}$, the set of atoms in the support by $\mathbb{S}_{+}=\{x\in\mathbb{S}:\pi(\{x\})>0\}$ and the set of non-atoms by $\mathbb{S}_{0}=\{x\in\mathbb{S}:\pi(\{x\})=0\}$. We define the inner product $\langle g|h\rangle_{\pi}=\int g(x)h(x)\pi(\mathrm{d}x)$ for $g,h\in L^{2}(\pi)$, the corresponding norm $\|f\|_{\pi}=\sqrt{\langle f|f\rangle_{\pi}}$, and the function spaces
@@
In our proofs, symmetricity of functions plays a central role, because of exchangeability of the variables.
@@
Let $n\in\mathbb{N}_{+}$ and $U\subset\mathbb{X}^{n}$. The function $f:U\to\mathbb{R}$ is symmetric if
@@
for any permutation $i_{1},\dots,i_{n}$ of $1,\dots,n$.
Let $N\in\mathbb{N}_{+}$, then in particular the mapping $z\mapsto\sum_{i=1}^{N}w(z_{i})$ is symmetric. We use the following notation for set of states where this mapping is positive:
@@ 6.1. Stationary covariances
because $w(z_{1})q(\mathrm{d}z_{1})=\pi(\mathrm{d}z_{1})$. Note also that $\langle f|P_{1}f\rangle=\langle f|f\rangle_{\pi}$. The following result shows that integrals similar to the one above can be written in a form involving a quadratic term.
@@
Let $N\in\mathbb{N}_{+}$, $M\in\mathbb{N}_{+}$, $f\in L^{2}(\pi)$ and $h:\mathbb{X}^{N+M}\to[0,\infty)$ be measurable bounded function and symmetric with respect to the first $N$ coordinates, then
@@
Since $h$ is symmetric with respect to the first $N$ coordinates and $z_{1},\dots,z_{N}$ are exchangeable under the integral, the expression on the left can be written as
@@
and if $B$ is $\pi$-positive we get similar inequality. ∎
@@
Let $N\in\mathbb{N}_{+}$, $f\in L^{2}(\pi)$ and let $h:\mathbb{X}^{N+2}\to[0,\infty)$ be measurable bounded function and symmetric with respect to the first $N+1$ coordinates, then
@@
If $\bar{f}\in L_{0,+}^{2}(\pi)$ and in addition for $z\in\mathbb{X}^{N+2}$ holds $h(z)>0$ whenever $z_{N+2}\in A$, where $A$ is a $\pi$-positive set, then the inequality is strict.
@@
Let us define $g_{n,k}:\mathbb{X}^{N+2}\to\mathbb{R}$,
$g_{n,k}(z_{1},\dots,z_{N+2})=\begin{cases}\frac{\big(\sum_{i=1,i\neq k}^{n}w(z_{i})f(z_{i})\big)^{2}}{\sum_{j=1,j\neq k}^{n}w(z_{j})},&\sum_{l=1,l\neq k}^{n}w(z_{l})>0\\ 0,&\sum_{l=1,l\neq k}^{n}w(z_{l})=0,\end{cases}$
for $n\in\{2,3,\dots,N+2\}$, $k\in\{1,2,\dots,n\}$. We have
(8) $\int_{\mathbb{X}^{N+2}}\frac{h(z)}{(N+1)N}\sum_{k=1}^{N+1}g_{N+1,k}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)\geq\int_{\mathbb{X}^{N+2}}\frac{h(z)}{N+1}g_{N+2,N+2}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
by Lemma 41 in Appendix C. The left hand side of (7) equals the left hand side of (8), since $h$ is symmetric with respect to the first $N+1$ coordinates and $z_{1},\dots,z_{N+1}$ are exchangeable under the integral.
If $\bar{f}\in L_{0,+}^{2}(\pi)$, then there exists two disjoint $\pi$-positive sets $U$ and $V$ such that $f$ differs on them, these sets and set AA are also $q$-positive. If we replace $\mathbb{X}^{N+2}$ with $U\times V\times\mathbb{S}^{N-1}\times A$ (for $N=1$ we omit $\mathbb{S}^{N-1}$) in the inequality (8) we get strict inequality by Lemma 41 and 42 in Appendix C, which implies that strict inequality also holds for (8). ∎
@@
$\langle f|P_{N}f\rangle_{\pi}\geq\langle f|P_{N+1}f\rangle_{\pi}.$
The inequality is strict if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
@@
$\langle f|P_{N}f\rangle-2\langle f|P_{N+1}f\rangle+\langle f|P_{N+2}f\rangle\geq 0.$
The inequality is strict if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
@@
(9) $\quad=2\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\sum_{i=1}^{N}w(z_{1})f(z_{1})w(z_{i})f(z_{i})}{\sum_{l=1}^{N}w(z_{l})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\qquad-2\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\sum_{i=1}^{N+1}w(z_{1})f(z_{1})w(z_{i})f(z_{i})}{\sum_{j=1}^{N+1}w(z_{j})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
@@
$\quad=\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\Big(\frac{B_{N+2}B_{N+1}A_{N}-2B_{N+2}B_{N}A_{N+1}+B_{N+1}B_{N}A_{N+2}}{B_{N+2}B_{N+1}B_{N}}\Big)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
@@
The claim follows, since $z_{N+1}$ and $z_{N+2}$ are exchangeable under the following integrals
@@
$\frac{2}{N}\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{l=1}^{N}w(z_{l})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\qquad-\frac{2}{N+1}\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\Big(\sum_{i=1}^{N+1}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N+1}w(z_{j})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
@@
Covariances can be expressed with the help of IID random variables from $q$. We will use this property and the fact that if $X\sim q$, then $\mathbb{E}[w(X)f(X)]=\mathbb{E}_{\pi}[f]$ for $\pi$-integrable $f$ and $\mathbb{E}[w(X)]=1$, to show the following limit property:
@@
Let $Z_{1},Z_{2},\dots$ be IID from $q$. Let us define sequences of random variables $(A_{N})$, $(B_{N})$, $(C_{N})$ and $(D_{N})$:
$A_{N}$ $=\frac{1}{N}\frac{\sum_{i=1}^{N}w(Z_{i})^{2}f(Z_{i})^{2}}{\sum_{j=1}w(Z_{j})},\qquad$ $B_{N}$ $=\frac{1}{N}\frac{\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}w(Z_{i})w(Z_{j})f(Z_{i})f(Z_{j})}{\sum_{k=1}^{N}w(Z_{k})},$
@@
By strong law of large numbers $C_{N}\to 0$ and $A_{N}+B_{N}\to\mathbb{E}_{\pi}[f]^{2}$ almost surely. Since $\mathbb{E}[C_{N}(Z)]\leq w(Z_{1})f(Z_{1})^{2}$ for all $N$, we have by dominated convergence $\mathbb{E}[C_{N}]\to 0$ and because $Z_{i}$ are exchangeable we have $\mathbb{E}[A_{N}]=\mathbb{E}[C_{N}]\to 0$, hence $A_{N}\to 0$ in probability and we obtain $B_{N}\to\mathbb{E}_{\pi}[f]^{2}$ in probability. It holds that $\mathbb{E}[|B_{N}|^{2}]\leq\mathbb{E}[D_{N}]$ and for $N\geq 2$ we have
@@
where the inequality follows from Sedrakyan’s inequality (Lemma 41 in Appendix C) and the equality from the fact that $Z_{1},\dots,Z_{N}$ are IID. Thus $\sup_{N}\mathbb{E}[|B_{N}|^{2}]<\infty$, which implies that $(B_{N})$ is uniformly integrable and therefore $\mathbb{E}[B_{N}]\to E_{\pi}[f]^{2}$. ∎
@@ 6.2. Rejection probabilities
Let $N\in\mathbb{N}_{+}$ then
$\epsilon(N)-\epsilon(N+1)>0,\qquad\epsilon(N)-2\epsilon(N+1)+\epsilon(N+2)>0$
and if $z_{1}\in\mathbb{S}$, then
$\epsilon(N,z_{1})-\epsilon(N+1,z_{1})>0,\qquad\epsilon(N,z_{1})-2\epsilon(N+1,z_{1})+\epsilon(N+2,z_{1})>0.$
@@
$\quad=w(z_{1})\int_{\mathbb{X}^{N+1}}\frac{1}{B_{N}}-\frac{2}{B_{N+1}}+\frac{1}{B_{N+2}}\bigg(\prod_{n=2}^{N+2}q(\mathrm{d}z_{n})\bigg)$
@@
where the third equality follows from Lemma 43 in Appendix C, and the fourth from
@@
Rejection probability and holding probability can be expressed with the help of IID random variables from $q$. We will use this property to show the following lemma:
@@
Let $Z_{2},Z_{3},\dots$ be IID from $q$, then
@@
In both equations the term inside the expectation is dominated by $1$ for all $N$ and by strong law of large numbers these terms converge to $\pi(\{x\})$ and 0 almost surely as $N\to\infty$. Therefore by dominated convergence we obtain the claim for holding probability and rejection probability. Thus, the claim follows for average rejection probability by dominated convergence. ∎

### v2

@@ 6. Convexity and monotonicity of i-SIR covariances and acceptance rates
We will now give the definitions we require. We denote the natural numbers as $\mathbb{N}=\{0,1,2,\dots\}$ and positive natural numbers as $\mathbb{N}_{+}=\{1,2,\dots\}$. Throughout the upcoming sections we have fixed $\pi$ and $q$ as mentioned in Section 1 (more detailed in Appendix A). We use the shorthand notation $\pi(\mathrm{d}x)=\pi(x)\mathrm{d}\mu(x)$ and $q(\mathrm{d}x)=q(x)\mathrm{d}\mu(x)$, where $\mu$ is the dominating measure. When we speak of atoms, we mean atoms with respect to $\pi$. The support of $\pi$ is denoted by $\mathbb{S}$, the set of atoms in the support by $\mathbb{S}_{+}=\{x\in\mathbb{S}:\pi(\{x\})>0\}$ and the set of non-atoms by $\mathbb{S}_{0}=\{x\in\mathbb{S}:\pi(\{x\})=0\}$. We define the inner product $\langle g|h\rangle_{\pi}=\int g(x)h(x)\pi(\mathrm{d}x)$ for $g,h\in L^{2}(\pi)$, the corresponding norm $\|f\|_{\pi}=\sqrt{\langle f|f\rangle_{\pi}}$, and the function spaces
@@
In our proofs, (permutation) symmetricity of functions plays a central role, because of exchangeability of the variables.
@@
Let $\in\mathbb{N}_{+}, m<nm<n$, $U\subset\mathbb{X}^{n}$ and $f:U\to\mathbb{R}$. The function is called symmetric if
@@
for any permutation $i_{1},\dots,i_{n}$ of $1,\dots,n$, and symmetric with respect to the first $m$ arguments if
$f(x_{1},\dots,x_{m},x_{m+1},\dots x_{n})=f(x_{i_{1}},\dots,x_{i_{m}},x_{m+1},\dots x_{n})$
for any permutation $i_{1},\dots,i_{m}$ of $1,\dots,m$.
Note that in Definition 11 the arguments $x_{i}\in\mathbb{X}$. Let $N\in\mathbb{N}_{+}$, then in particular the mapping $z\mapsto\sum_{i=1}^{N}w(z_{i})$ is symmetric. We use the following notation for set of states where this mapping is positive:
@@ 6.1. Stationary covariances
because $w(z_{1})q(\mathrm{d}z_{1})=\pi(\mathrm{d}z_{1})$. Note also that $\langle f|P_{1}f\rangle=\langle f|f\rangle_{\pi}$. The following result shows that integrals similar to the one above can be written in a form involving a quadratic term:
@@
Let $N,M\in\mathbb{N}_{+}$, $f\in L^{2}(\pi)$ and $h:\mathbb{W}_{N}\times M\to[0,\infty)$ be measurable bounded function and symmetric with respect to the first $N$ arguments, then
@@
Since $h$ is symmetric with respect to the first $N$ arguments and $z_{1},\dots,z_{N}$ are exchangeable under the integral, the expression on the left can be written as
@@
and if $B$ is $\pi$-positive, we get similar inequality. ∎
@@
Let $N\in\mathbb{N}_{+}$, $f\in L^{2}(\pi)$ and let $h:\mathbb{W}_{N+1}\times\mathbb{X}\to[0,\infty)$ be measurable bounded function and symmetric with respect to the first $N+1$ arguments, then
@@
If $\bar{f}\in L_{0,+}^{2}(\pi)$ and there exists a $\pi$-positive set $A$ such that $h(z)>0$ for every $z\in\mathbb{W}_{N+1}\times A$, then the inequality is strict.
@@
Let us define $g_{k}:\mathbb{W}_{N+1}\times\mathbb{X}\to\mathbb{R}$,
$g_{k}(z_{1},\dots,z_{N+2})=\begin{cases}\frac{\big(\sum_{i=1,i\neq k}^{N+1}w(z_{i})f(z_{i})\big)^{2}}{\sum_{j=1,j\neq k}^{N+1}w(z_{j})},&\sum_{l=1,l\neq k}^{N+1}w(z_{l})>0\\ 0,&\sum_{l=1,l\neq k}^{N+1}w(z_{l})=0,\end{cases}$
for $k\in\{1,2,\dots,N+2\}$. We have
(8) $\int_{B}\frac{h(z)}{(N+1)N}\sum_{k=1}^{N+1}g_{k}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)\geq\int_{B}\frac{h(z)}{N+1}g_{N+2}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
for any $B\subset\mathbb{W}_{N+1}\times\mathbb{X}$ that is measurable with respect to the product sigma-algebra by Lemma 41 in Appendix C. When $B=\mathbb{W}_{N+1}\times\mathbb{X}$, the right-hand side of (8) equals the right-hand side of (7), and
$\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{h(z)}{(N+1)N}\sum_{k=1}^{N+1}g_{k}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)=\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{h(z)}{N}g_{N+1}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{h(z)}{N}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg),$
where the first equality follows since $h$ is symmetric with respect to the first $N+1$ arguments and $z_{1},\dots,z_{N+1}$ are exchangeable under the integral, and the second from $\mathbb{W}_{N}\times\mathbb{X}\subset\mathbb{W}_{N+1}$.
If $\bar{f}\in L_{0,+}^{2}(\pi)$, then there exist two disjoint $\pi$-positive sets $U,V\subset\mathbb{S}$ such that $f$ differs on them, these sets and set AA in the assumption are $q$-positive. Let $W=U\times V\times\mathbb{S}^{N-1}\times A$ (for $N=1$, $W=U\times V\times A$). If $B=W$, we get strict inequality for (8) by Lemma 41 and 42 in Appendix C, thus strict inequality also holds when $B=\mathbb{W}_{N+1}\times\mathbb{X}=W\cup((\mathbb{W}_{N+1}\times\mathbb{X})\setminus W)$, from which the claim follows. ∎
@@
$\langle f|P_{N}f\rangle_{\pi}\geq\langle f|P_{N+1}f\rangle_{\pi},$
where the inequality is strict if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
@@
$\langle f|P_{N}f\rangle-2\langle f|P_{N+1}f\rangle+\langle f|P_{N+2}f\rangle\geq 0,$
where the inequality is strict if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
@@
(9) $\quad=2\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\sum_{i=1}^{N}w(z_{1})f(z_{1})w(z_{i})f(z_{i})}{\sum_{k=1}^{N}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\qquad-2\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\sum_{i=1}^{N+1}w(z_{1})f(z_{1})w(z_{i})f(z_{i})}{\sum_{k=1}^{N+1}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
@@
$\quad=\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\frac{B_{N+2}B_{N+1}A_{N}-2B_{N+2}B_{N}A_{N+1}+B_{N+1}B_{N}A_{N+2}}{B_{N+2}B_{N+1}B_{N}}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
@@
The claim follows since $z_{N+1}$ and $z_{N+2}$ are exchangeable under the following integrals
@@
$\frac{2}{N}\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{k=1}^{N}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\qquad-\frac{2}{N+1}\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\Big(\sum_{i=1}^{N+1}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{k=1}^{N+1}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
@@
Covariances can be expressed with the help of i.i.d. random variables from $q$. We will use this property and the fact that if $X\sim q$, then $\mathbb{E}[w(X)f(X)]=\mathbb{E}_{\pi}[f]$ for $\pi$-integrable $f$ and $\mathbb{E}[w(X)]=1$, to show the following limit property:
@@
Let $Z_{1},Z_{2},\dots$ be i.i.d. from $q$. Let us define sequences of random variables $(A_{N})$, $(B_{N})$, $(C_{N})$ and $(D_{N})$:
$A_{N}$ $=\frac{1}{N}\frac{\sum_{i=1}^{N}w(Z_{i})^{2}f(Z_{i})^{2}}{\sum_{j=1}^{N}w(Z_{j})},\qquad$ $B_{N}$ $=\frac{1}{N}\frac{\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}w(Z_{i})w(Z_{j})f(Z_{i})f(Z_{j})}{\sum_{k=1}^{N}w(Z_{k})},$
@@
By strong law of large numbers $C_{N}\to 0$ and $A_{N}+B_{N}\to\mathbb{E}_{\pi}[f]^{2}$ almost surely. Since $|C_{N}(Z)|\leq w(Z_{1})f(Z_{1})^{2}$ for all $N$, we have by dominated convergence $\mathbb{E}[C_{N}]\to 0$ and because $Z_{i}$ are exchangeable, we have $\mathbb{E}[A_{N}]=\mathbb{E}[C_{N}]\to 0$, hence $A_{N}\to 0$ in probability and we obtain $B_{N}\to\mathbb{E}_{\pi}[f]^{2}$ in probability. It holds that $\mathbb{E}[|B_{N}|^{2}]\leq\mathbb{E}[D_{N}]$ and for $N\geq 2$ we have
@@
where the inequality follows from Sedrakyan’s inequality (Lemma 40 in Appendix C) and the equality from the fact that $Z_{1},\dots,Z_{N}$ are i.i.d. Thus, $\sup_{N}\mathbb{E}[|B_{N}|^{2}]<\infty$, which implies that $(B_{N})$ is uniformly integrable and therefore $\mathbb{E}[B_{N}]\to E_{\pi}[f]^{2}$. ∎
@@ 6.2. Rejection probabilities
Let $N\in\mathbb{N}_{+}$ and $x\in\mathbb{S}$, then
$\epsilon(N)-\epsilon(N+1)$ $>0,\qquad$ $\epsilon(N)-2\epsilon(N+1)+\epsilon(N+2)$ $>0,$
$\epsilon(N,x)-\epsilon(N+1,x)$ $>0,\qquad$ $\epsilon(N,x)-2\epsilon(N+1,x)+\epsilon(N+2,x)$ $>0.$
@@
where the second equality follows from Lemma 43 in Appendix C, and the third from
@@
Rejection probability and holding probability can be expressed with the help of i.i.d. random variables from $q$. We will use this property to show the following lemma:
@@
Let $Z_{2},Z_{3},\dots$ be i.i.d. from $q$, then
@@
In both equations the term inside the expectation is dominated by $1$ for all $N$ and by strong law of large numbers the term in the first equation converges to $\pi(\{x\})$ and term in the second equation converges to 0 almost surely as $N\to\infty$. Therefore, by dominated convergence we obtain the claim for holding probability and rejection probability. Thus, the claim for average rejection probability follows by dominated convergence. ∎